What volumes of 60% hydrocloric acid solution and 30% hydrocloric acid solution must be mixed to make 125 ml of 36% hydrocloric acid solution.
Answer by checkley77(12844)

.60X+.30(125-X)=.36*125
.60X+37.5-.30X=45
.30X=45-37.5
.30X=7.5
X=7.5/.3
X=25 ML. OX 60% ACID IS USED.
125-25=100 ML. OF 30% ACID IS USED.
.60*25+.30*100=.368125
15+30=45
45=45